# Complements of discriminants of real parabolic function singularities. II

### v3

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
Proof. The restriction of the polynomial (61) to the plane $\{l=0\}$ is a non-degenerate homogeneous cubic polynomial that vanishes on three real lines, see Fig. 10 (left). This implies the first assertion of the lemma analogously to the proof of Lemma 39. Let $a$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$. The restriction of this polynomial to the line connecting 0 and $a$ is a polynomial of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0. $\Box$
Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin. These numbers form a lower semi-continuous function on $L$. By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$. The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant. This union is connected and invariant under the action (44) of the group $S(3)$. By Lemma 47, its point corresponding to $l\equiv x-y$ belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$. Therefore, this entire component is $S(3)$-invariant. This concludes the proof of Theorem 42 for the virtual component with $\mbox{Card}=1216$.
@@
Consider a tubular neighborhood of this stratum in $\Xi_{2}$ fibered into its orthogonal slices diffeomorphic to small four-dimensional balls centered at the points of the stratum. Suppose it is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is a versal deformation of the critical point of the function $F(l)$ at the origin. Therefore, the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
@@
Denote the fiber over the point $l\in L$ by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
@@
Similarly, define the subset $W(l)\subset U(l)$ as the union of points satisfying condition (a) and condition (b) replaced by $\mbox{Card}(f)=1318$.
@@
To prove Lemma 49, we will describe detail the geometry of the discriminant in the parameter space of the versal deformation (62) in detail. The Morse polynomials of the form (62) can have two or four critical points depending on the parameters $\varkappa_{2},\varkappa_{3},$ and $\varkappa_{4}$. In the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse functions looks as shown in Fig. 11 (left), see [3], § 21.3.
@@
This set divides the space ${\mathbb{R}}^{3}$ into three parts. The larger part consists of polynomials with only two real critical points. Two symmetric “pyramids” are filled with polynomials (62) that have three saddlepoints and one point that is either a minimum or a maximum, depending on the pyramid. We are only interested in the polynomials (62) with a minimum point. For each generic interior point $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ of the corresponding pyramid, the set of non-discriminant polynomials (62) with these values consists of five intervals. Indeed, the coefficient $\varkappa_{1}$ of a polynomial from this set should be not equal to minus a critical value of the polynomial
@@
i.e., it can belong to one of the five intervals into which the critical values divide the line. For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal. The set of these non-generic polynomials (63) divides the pyramid into six parts. The intersection of the pyramid with a plane perpendicular to its axis looks as the curvilinear triangle shown in Fig. 11 (center). The non-generic polynomials form three segments in it. In particular, the central point of this triangle corresponds to a polynomial $f_{\triangle}$ with equal critical values at all three saddlepoints. The zero-level set of the function $f_{\triangle}-\tau$, where $\tau$ is the critical value, looks as in Fig. 10 (right). These polynomials, representing various sections of the pyramid, sweep out its central axis. An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
@@
Let us label the three saddlepoints of an arbitrary polynomial inside the pyramid as $a,b,c$. This marking of critical points can be uniquely extended by continuity to all other polynomials within the pyramid. Each of the six cells into which the curvilinear triangle in Fig. 11 (center) is split by the segments is characterized by the order in $\mathbb{R}$} of the critical values of the corresponding polynomials at the marked points.
@@
The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of nine connected domains. We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value. The projections of these domains to the space ${\mathbb{R}}^{3}$ of coefficients $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ are three shares of the pyramid: their intersections with a section of the pyramid are shown in Fig. 11 (right). Each share, marked by a letter $a,b,$ or $c$ in the picture, consists of polynomials (63) whose critical value at the saddlepoint with that marking is the highest. The corresponding domain in the parameter space ${\mathbb{R}}^{4}$ of the family (62) is a fiber bundle over the corresponding share shown in Fig. 11 (right), where the fiber over a polynomial (63) is equal to the interval between two its highest critical values. Clearly, all three of these domains are contractible.
@@
Proof of Lemma 49. For each $l\in L$, the orthogonal slice $U(l)$ of the discriminant stratum $\{D_{4}^{-}\}\subset\Xi_{2}$ at the point $\{F(l)\}$ is a versal deformation of the corresponding singularity of class $D_{4}^{-}$. The inducing map from the definition of versal deformations (see [3], § 8) identifies this slice with a neighborhood of the origin in the parameter space ${\mathbb{R}}^{4}$ of the deformation (62) and maps the discriminant to the discriminant. Consider three domains in the slice $U(l)$ identified by this map with the three domains described in the previous paragraph. These three domains in $U(l)$ contain all the points in the slice that correspond to polynomials having exactly three critical points with negative critical values: one minimum and two points with Morse index 1.
@@
Consider again the case $l\equiv x-y$. Let $f$ be a point in one of these three domains in the slice $U(l)$ at the point $F(l)$. We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\zeta$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points. Then, acting as in the proof of Lemma 47, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\zeta$ and $\tilde{\omega}_{1}$. For the same reasons as in the aforementioned proof, the corresponding critical points meet at a real critical point of class $A_{2}$. Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points. We again obtain a polynomial with Coxeter-Dynkin graph shown in Fig. 7. This polynomial has four critical points with Morse index 1; three of these points arise from the splitting of the $D_{4}^{-}$ singularity. The three corresponding vertices in the middle row of this graph are the leaf vertex and some two of the vertices connected to upper vertices. Only one of these three critical points has a positive critical value. Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex. This choice explicitly determines the virtual component associated with the obtained polynomial (and? consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement). Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant. It turns out that $\mbox{Card}=844$ when the right-hand vertex in Fig. 7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other cases. $\Box$
@@
As $l$ moves along half of the circle $L$, the polynomial $F(l)$ and the corresponding slice $U(l)$ returns to themselves. The union of the contractible domains $V(l)\subset U(l)$ over all $l$ in this half-circle is a connected set that is $S(3)$-invariant. All of these domains consist of polynomials with $\mbox{Card}=844$. This proves the assertion of Theorem 42 for the virtual components with $\mbox{Card}=844$.
@@
The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ of all polynomials $F(l)$ permutes the two components of the sets $W(l)$ mentioned in Lemma 49.
@@
Figure 12. The plane $\{l=0\}$ and the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
@@
(64) $x^{3}+y^{3}+z^{3}-6xyz=0,$
which is homeomorphic to the quadratic cone in ${\mathbb{R}}^{3}$. The third line belongs to the component that is homeomorphic to the plane, see Fig. 12 (left).
Consider the local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$ at the origin in ${\mathbb{R}}^{3}$ such that $l\approx\tilde{l}$ up to terms of second degree, and $F(l)$ has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class. The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin. The curves on this surface that are distinguished by the conditions $\tilde{p}=0$, $\tilde{p}=\tilde{q}$, and $\tilde{p}=-\tilde{q}$ are tangent to the three lines considered in the previous paragraph. Let $\hat{F}(l)\in U(l)$ be a perturbation of $F(l)$ that splits the $D_{4}^{-}$ singularity into a minimum point with critical value $\varepsilon$ and three critical points of Morse index 1 with critical value 0. In canonical coordinates, the zero-level set of the restriction of $\hat{F}(l)$ to the surface $\{\tilde{l}=0\}$ locally consists of three curves obtained by small deformations of the three curves for the function $F(l)$, and the three critical points occur at their intersections, see Fig. 10 (right). Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the nearby domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at exactly this critical point is positive and the critical values at the other two such points are negative. Thus, there is a canonical one-to-one correspondence between these domains in $U(l)$ and the choices of some two out of three lines in the plane $\ker l$ considered in the previous paragraph.
As $l$ moves along the half-circle $L/{\mathbb{Z}}_{2}$, these three lines also move in such a way that one pair of them returns to itself, while the other two pairs permute. Thus, the monodromy over this half-circle acts in the same way on the set of three related domains in $U(l)$. We know one of these domains which is moved by this monodromy to itself: it is the domain $V(l)$ considered in Lemma 49. Therefore, the permuted domains belong to the set $W(l)$. $\Box$
@@
Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary linear function of unit norm, whose kernel is tangent to the component of the surface ((64)) that is homeomorphic to the quadratic cone, see Fig. 12 ((right)). The polynomial
@@
then has a singularity of class $D_{5}^{-}$ at the origin in ${\mathbb{R}}^{3}$. It has the form
@@
Proof. All singularities of corank 2 that can occur in the deformation (42) of $P_{8}^{2}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [19], [12]. The restriction of the polynomial ((65)) to the plane $\ker l$ is a degenerate cubic form; therefore, its singularity at the origin is more complicated than $D_{4}$. According to Corollary 10.4.9 of [21], this singularity is not of class $E_{6}$. Thus, it remains to choose between the classes $D_{5}^{+}$ and $D_{5}^{-}$. In the particular case when
@@
the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and Y=y3zY=y-3z reduce the polynomial (65) to a polynomial whose lower quasihomogeneous part is equal to $\tilde{l}^{2}+\frac{63}{2}Y^{2}z-\frac{9}{4}z^{4}$. A dilation of coordinates reduces it to the form (66).
@@
For any linear function $l$ as in Lemma 51, the set of polynomials with $\mbox{Ind}=-3$ consists of a single connected component in a small neighborhood of the corresponding polynomial ((65)).
@@
  • the critical values at all points with Morse index 1 are lower than all critical values at the points with Morse index 2,
@@
Indeed, this polynomial can be obtained from the perturbation of (65) shown in the middle part of Fig. 13 by a minor additional perturbation that properly orders the critical values. The intersection indices of its critical values are derived using the Gusein-Zade—A’Campo method [10].
Our program shows that only six different virtual functions of $P_{8}^{2}$ class satisfy these conditions. It also shows that all six of these virtual functions appear in the virtual component with $\mbox{Card}=156.$ $\Box$
The set of all polynomials (65) corresponding to different tangent planes of the cone is connected and $S(3)$-invariant. The same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}\}$ stratum at the these points. This implies the statement of Theorem 42 for the virtual component with $\mbox{Card}=156$. $\Box$

### v4

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
Proof. The restriction of the polynomial (61) to the plane $\{l=0\}$ is a non-degenerate homogeneous cubic polynomial that vanishes on three real lines, see Fig. 10 (left). The first assertion of the lemma follows from this analogously to the proof of Lemma 39. Let $a$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$. The restriction of the polynomial to the line through 0 and $a$ is a polynomial function of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0. $\Box$
Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin. These numbers $c(l)$ form a lower semi-continuous function on $L$. By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$. The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant. This union is connected and invariant under the action (44) of the group $S(3)$. By Lemma 47, for $l\equiv x-y$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$. The polynomial $F(l)-\delta$ also belongs to this component because all the polynomials F(l)tF(l)-t, where $t\in[\delta,1]$, are non-discriminant. Therefore, the entire component is $S(3)$-invariant. This concludes the proof of Theorem 42 for the virtual component with $\mbox{Card}=1216$.
@@
Consider a tubular neighborhood of this stratum in $\Xi_{2}$ fibered into its orthogonal slices diffeomorphic to small four-dimensional balls centered at the points of the stratum. Suppose it is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is a versal deformation of the critical point of the function $F(l)$ at the origin. Therefore, all the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
@@
Denote the fiber of this tubular neighborhood over the point $l\in L$ by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
@@
Similarly, define the subset $W(l)\subset U(l)$ as the union of points satisfying condition (a) but with condition (b) replaced by $\mbox{Card}(f)=1318$.
@@
To prove Lemma 49, we will describe the geometry of the discriminant in the parameter space of the versal deformation (62) in detail. The Morse polynomials of the form (62) can have two or four critical points depending on the parameters $\varkappa_{2},\varkappa_{3},$ and $\varkappa_{4}$. In a neighborhood of the origin of the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse functions looks as shown in Fig. 11 (left), see [3], § 21.3.
@@
This set divides the space ${\mathbb{R}}^{3}$ into three parts. The larger part consists of polynomials with only two real critical points. Two symmetric “pyramids” are filled with polynomials (62) that have three saddlepoints and one critical point that is either a minimum or a maximum, depending on the pyramid. We are only interested in the polynomials (62) with a minimum point. For each generic interior point $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ of the corresponding pyramid, the set of non-discriminant polynomials (62) with these values consists of five intervals. Indeed, the coefficient $\varkappa_{1}$ of a polynomial from this set should be not equal to minus a critical value of the polynomial
@@
i.e., it can belong to one of the five intervals into which the critical values divide the line $\mathbb{R}$}. For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal. The set of these non-generic polynomials (63) divides the pyramid into six parts. The intersection of the pyramid with a plane perpendicular to its axis looks as the curvilinear triangle shown in Fig. 11 (center). The non-generic polynomials form three segments in it. In particular, the central point of this triangle corresponds to a polynomial $f_{\triangle}$ with equal critical values at all three saddlepoints. The zero-level set of the function $f_{\triangle}-\tau$, where $\tau$ is the critical value, looks as in Fig. 10 (right). These polynomials, representing various sections of the pyramid and corresponding to different values of $\tau$, sweep out its central axis. An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
@@
Let us label the three saddlepoints of an arbitrary polynomial inside the pyramid as $a,b,c$. This labeling of critical points can be uniquely extended by continuity to all other polynomials within the pyramid. Each of the six cells, into which the curvilinear triangle in Fig. 11 (center) is split by the segments, is characterized by the order in $\mathbb{R}$} of the critical values of the corresponding polynomials at the marked points.
@@
The set of non-discriminant polynomials (62) whose parameters $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ belong to this pyramid consists of eight connected domains. We are only interested in three of these domains, whose points are the polynomials (62) which have exactly one positive critical value. The projections of these domains to the space ${\mathbb{R}}^{3}$ of coefficients $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ are three shares of the pyramid: their intersections with a section of the pyramid are shown in Fig. 11 (right). Each share, marked by a letter $a,b,$ or $c$ in the picture, consists of polynomials (63) whose critical value at the saddlepoint with that marking is the highest. The corresponding domain in the parameter space ${\mathbb{R}}^{4}$ of the family (62) is a fiber bundle over the corresponding share shown in Fig. 11 (right), where the fiber over a polynomial (63) is equal to the interval between two its highest critical values. Clearly, all three of these domains are contractible.
@@
Proof of Lemma 49. For each $l\in L$, the orthogonal slice $U(l)$ of the discriminant stratum $\{D_{4}^{-}\}\subset\Xi_{2}$ at the point $\{F(l)\}$ is a versal deformation of the corresponding singularity of class $D_{4}^{-}$. The inducing map from the definition of versal deformations (see [3], § 8) identifies this slice with a neighborhood of the origin in the parameter space ${\mathbb{R}}^{4}$ of the deformation (62) and maps the discriminant to the discriminant. Consider three domains in the slice $U(l)$ identified by this map with the three domains described in the previous paragraph. These three domains in $U(l)$ contain all the points in the slice that correspond to polynomials of class $\Xi_{2}$ having exactly three critical points with negative critical values: one minimum and two points with Morse index 1.
@@
Consider again the case $l\equiv x-y$. Let $f$ be a point in one of these three domains in the slice $U(l)$ at the point $F(l)$. We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\zeta$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points. Then, acting as in the proof of Lemma 47, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\zeta$ and $\tilde{\omega}_{1}$. For the same reasons as in the proof of Lemma 47, the corresponding critical points meet at a real critical point of class $A_{2}$. Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points. We again obtain a polynomial with Coxeter-Dynkin graph shown in Fig. 7. This polynomial has four critical points with Morse index 1; three of these points arise from the splitting of the $D_{4}^{-}$ singularity. The three corresponding vertices in the middle row of this graph are the leaf vertex and some two of the vertices connected to upper vertices. Only one of these three critical points has a positive critical value. Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex. This choice explicitly determines the virtual component associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement). Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant. It turns out that $\mbox{Card}=844$ when the right-hand vertex in Fig. 7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other cases. $\Box$
The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l).
@@
The union of the contractible domains $V(l)\subset U(l)$ over all $l$ in this circle is a connected set that is $S(3)$-invariant. All of these domains consist of polynomials with $\mbox{Card}=844$. This proves the assertion of Theorem 42 for the virtual components with $\mbox{Card}=844$.
The union of all sets $W(l)$ also is $S(3)$-invariant. It is the space of a fiber bundle over the same circle, with the fiber $W(l)$ consisting of two contractible domains.
@@
The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $D_{4}^{-}$ stratum at the corresponding polynomials $F(l)$.
@@
Figure 12. The plane sections of the surface $x^{3}+y^{3}+z^{3}-6xyz=0$
@@
(64) $x^{3}+y^{3}+z^{3}-6xyz=0$
that is homeomorphic to the quadratic cone in ${\mathbb{R}}^{3}$. The third line belongs to the component that is homeomorphic to the plane, see Fig. 12 left (in which we schematically depicted the set (64) by the union of the quadratic cone and the horizontal plane).
The singularity of the function $F(l)$ at the origin has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class in some local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$, where $\tilde{l}\approx l$ up to terms of second degree. The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin. The curves on this surface that are distinguished by the conditions
(65) $\tilde{p}=0,\ \tilde{p}=\tilde{q},\ \mbox{and}\ \tilde{p}=-\tilde{q}$
are tangent to the three lines considered in the previous paragraph. Consider the set of points in the four-dimensional slice $U(l)$, which correspond to polynomials, at which the $D_{4}^{-}$ singularity of $F(l)$ splits into a minimum point with negative critical value and three critical points with Morse index 1 and critical value 0. This set is locally homeomorphic to a single interval with an endpoint at the point $\{F(l)\}$: it corresponds to the axis of one of pyramids in Fig. 11 left. Each function $\hat{F}(l)$ from this set has the form $\check{l}^{2}+\varphi(\check{p},\check{q})$ in local coordinates $\check{l},\check{p},\check{q}$, which are very close to $\tilde{l},\tilde{p},\tilde{q}$, where the zero-level set of the function $\varphi$ looks as in Fig. 10 right. The three curves constituting this zero-level set in the surface $\{\check{l}=0\}$ are very close to three curves (65) in the set $\{\tilde{l}=0\}$. In particular, they are in a canonical one-to-one correspondence with these curves, and hence also with three lines of the intersection of the surface (64) and the plane $\ker l$. The three critical points with Morse index 1 occur at their intersections, see Fig. 10 (right). Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at the critical point obtained by this perturbation from the chosen critical point of $\hat{F}(l)$ is positive and the critical values at the other two critical points with Morse index 1 are negative. Thus, there is a canonical one-to-one correspondence between these three domains in $U(l)$ and the choices of some two out of three lines distinguished by the equations (64) and $l=0$.
As $l$ moves along the circle $L/{\mathbb{Z}}_{2}$, these three lines also move in such a way that one pair of them returns to itself, while the other two pairs permute. Thus, the monodromy over this circle acts in the same way on the set of three related domains in $U(l)$. We know one of these domains which is moved by this monodromy to itself: it is the domain $V(l)$ considered in Lemma 49. Therefore, the permuted domains belong to the set $W(l)$. $\Box$
@@
Let $l:{\mathbb{R}}^{3}\to{\mathbb{R}}$ be an arbitrary linear function of unit norm, whose kernel is tangent to the component of the surface ((64)) that is homeomorphic to the quadratic cone, see Fig. 12 right. The polynomial
@@
then has a singularity of class $D_{5}^{-}$ at the origin in ${\mathbb{R}}^{3}$. In particular, it has the form
@@
Proof. All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [20], [12]. The restriction of the polynomial ((66)) to the plane $\ker l$ is a degenerate cubic form; therefore, its singularity at the origin is more complicated than $D_{4}$. According to Corollary 10.4.9 of [22], this singularity is not of class $E_{6}$. Thus, it remains to choose between the classes $D_{5}^{+}$ and $D_{5}^{-}$. In the particular case when
@@
the substitutions $\tilde{l}=l+\left(\frac{3}{16}(5y-11z)^{2}-\frac{3}{2}yz\right)$ and $\tilde{y}=y-3z$ reduce the polynomial (66) to a polynomial whose lower weighted homogeneous part is equal to $\tilde{l}^{2}+\frac{63}{2}\tilde{y}^{2}z-\frac{9}{4}z^{4}$. A dilation of coordinates reduces it to the form (67), and technique of [3], §12, removes all higher weighted homogeneous terms of the Taylor expansion.
@@
For any linear function $l$ as in Lemma 51, the set of polynomials with $\mbox{Ind}=-3$ in a small neighborhood of the corresponding polynomial ((66)) consists of a single connected component.
@@
  • the critical values at all critical points with Morse index 1 are lower than all critical values at the points with Morse index 2,
@@
Indeed, a polynomial with these properties can be obtained from the perturbation of (66) shown in the middle part of Fig. 13 by a minor additional perturbation that properly orders the critical values. The intersection indices of its vanishing cycles are derived using the Gusein-Zade—A’Campo method [10].
Our program, upon a special request, lists all virtual functions of $P_{8}^{2}$ class satisfying these conditions. It turns out that there are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$
The set of all polynomials (66) corresponding to different tangent planes of the cone is connected and $S(3)$-invariant. According to Lemma 52, the same is true for the union of all local domains with $\mbox{Card}=156$ in the transversal slices to the $\{D_{5}^{-}\}$ stratum at these points. Indeed, this union is the space of a fiber bundle with contractible fibers over the space of such planes. This implies the statement of Theorem 42 for the virtual component with $\mbox{Card}=156$. $\Box$